Construction Cost Analysis and Estimating:

A project requires 17,000 allowed hours. Location of the work is above the arctic circle, and the efficiency is estimated as 45%. Is the gross hourly cost is $50, then what is the estimated cost? If the work is in Texas, hourly cost is $30, and an efficiency of 85% is assumed, then what is the cost of direct labor?
Answer #1

Hours required = 17,000

Arctic Circle:-

Efficiency = 45%

Total working hours required= 17,000/0.45= 37,770

Grossly hour rate = $50

Total cost of direct labour = 37,770*50=$188,500

Texas

Hourly cost = $30

Efficiency= 85%

Total working hours required= 17,000/0.85= 20,000

Direct labour cost= 20,000*30=$60,0000